### v2

# Antenna Coding Empowered by Pixel Antennas

Abstract
Pixel antennas, based on discretizing a continuous radiation surface into small elements called pixels, are a flexible reconfigurable antenna technology. By controlling the connections between pixels via switches, the characteristics of pixel antennas can be adjusted to enhance the wireless channel. Inspired by this, we propose a novel technique denoted antenna coding empowered by pixel antennas. We first derive a physical and electromagnetic based communication model for pixel antennas using microwave multiport network theory and beamspace channel representation. With the model, we optimize the antenna coding to maximize the channel gain in a single-input single-output (SISO) pixel antenna system and develop a codebook design for antenna coding to reduce the computational complexity. We analyze the average channel gain of SISO pixel antenna system and derive the corresponding upper bound. In addition, we jointly optimize the antenna coding and transmit signal covariance matrix to maximize the channel capacity in a multiple-input multiple-output (MIMO) pixel antenna system. Simulation results show that using pixel antennas can enhance the average channel gain by up to 5.4 times and channel capacity by up to 3.1 times, demonstrating the significant potential of pixel antennas as a new dimension to design and optimize wireless communication systems.
I Introduction
Antennas play a significant role in wireless communication systems. Utilizing multiple antennas, multiple-input multiple-output (MIMO) systems have been one of the most celebrated wireless communication technologies over the past few decades. It has evolved from point-to-point MIMO in the third generation (3G) to multiuser MIMO in the fourth generation (4G) and massive MIMO in the fifth generation (5G), and will still be a dominating technology in the upcoming sixth generation (6G) [1]. However, the antennas utilized in conventional MIMO technologies have fixed configurations and characteristics such as operating frequency, radiation pattern, and polarization, which are not involved in the signal processing and system optimization in wireless communications. As 6G mobile communication strives to push the performance to extreme levels far beyond current 5G mobile communications, it is necessary to seek a paradigm shift in antenna technology to achieve a new degree of freedom in designing and optimizing wireless communication systems, so as to break through the current performance limit.
Pixel antennas are a flexible antenna technology to design highly reconfigurable antennas and dates back to 2004 [2], [3]. The concept of the pixel antenna is based on discretizing a continuous radiation surface into small elements, which are referred to as pixels, and connecting adjacent pixels through hardwires or RF switches [4, 5, 6], [7], [8], [9] and [10]. By controlling the connections between pixels, various antenna topologies can be formed so that the pixel antenna can be reconfigured to achieve a wide range of distinct antenna characteristics such as operating frequency, radiation pattern, and polarization. For example, a frequency-reconfigurable pixel antenna, which can reconfigure the operating frequency from PCS band to UMTS band, has been designed in [7], and pattern-reconfigurable pixel antennas, which can reconfigure the radiation pattern to implement $360\text{\textdegree}$ single- or multi-beam steering, have been designed in [8], [9], and [10]. Moreover, various pixel antennas have been designed for different applications. In [11], a reconfigurable intelligent surface with pixelated elements, which can reconfigure the phase shift of reflected waves, has been designed. In [12], and [13], multiport pixel rectennas, which can maximize the output dc power for ambient RF energy harvesting, have been developed. In [14], MIMO antennas with pixelated surface, which provide low mutual coupling and spatial correlation to maximize the ergodic capacity, have been developed. In [15], defected ground structures with pixelated grid, which suppress cross-polarization and achieve circular polarization, have been designed. While these works [3], [4, 5, 6], [7], [8], [9], [10], [11], [12], [13], [14], [15] have designed pixel antennas with significant performance enhancement, the investigations are limited at the level of antenna hardware design and do not consider the impact on wireless communications at the level of the system. Thus, it is necessary to explore using pixel antennas to design and optimize wireless communication systems.
In parallel with the development of pixel antennas, the concept of the fluid antenna system (FAS) has also been proposed and introduced in 2020 [16]. FAS originated from the emergence of mechanically flexible antennas such as liquid antennas based on liquid metals or ionized solutions [17], which enables a single antenna freely switching the position in a small linear space. Moreover, FAS can be implemented by any appropriate reconfigurable technique including electronically flexible pixel antennas, where the pixels can be turned on-and-off instantly so that several pixels can be turned on to form an antenna port in a particular position. One of the first pixel antenna designs for an FAS has been demonstrated in [18] to show that a single pixel antenna can mimic switching positions in a small linear space and be useful for FAS. Leveraging the position-switchable property, FAS can select the optimal position in a given space to adapt to fading channels [19] and thus enhance the system performance such as level crossing rate [16], diversity gain [20], and outage probability [21]. FAS has also been utilized to assist MIMO communication systems [22], [23], [24], [25]. Specifically, the capacity, diversity-multiplexing tradeoff, spectral efficiency, and energy efficiency for MIMO-FAS have been investigated in [22], [23], [24] and the rate maximization for MIMO-FAS based on statistical channel state information (CSI) has been studied in [25]. In addition, a novel multiple access based on FAS, denoted as fluid antenna multiple access (FAMA), has been proposed in [26], and FAMA can be categorized into fast FAMA [27] and slow FAMA [28], Moreover, FAS has been investigated to enhance emerging areas such as secret communications [29] and full duplex communications [30].
While the emergence and development of FAS [16], [19], [20], [21], [22], [23], [24], [25], [26], [27], [28], [29], [30] has preliminarily demonstrated the potential of pixel antennas for designing and optimizing wireless communication system, there is much further potential for pixel antennas to be exploited in wireless communication. For example, utilizing the individual pixels inside a pixel antenna, and the connections between them, can open up new possibilities for their use in wireless communication. To fully exploit the potential of pixel antennas for wireless communication, there remains two open problems: 1) how to derive a physical and electromagnetic (EM) based communication model for pixel antennas, and 2) how to efficiently utilize the flexibility of pixel antennas to enhance wireless communications.
In this paper, we propose using pixel antennas to enhance wireless communication systems with an EM based model and a systematic and efficient approach. In comparison with conventional wireless communication systems using antennas with fixed configuration, using pixel antennas allows designing and optimizing wireless communication systems in a new dimension through adjusting the antenna configuration, and thus the wireless communication systems performance can be significantly enhanced. The contributions of the paper are summarized as follows.
First, we propose a novel technique called antenna coding empowered by pixel antennas to enhance wireless communication systems. By adjusting the connections between adjacent pixels through switches, the characteristics of the pixel antenna such as impedance, radiation pattern, and polarization can be controlled. Inspired by this, we propose antenna coding, which is a technique to control antenna configuration and characteristics through binary codes representing switches for enhancing wireless systems. With antenna coding, wireless systems can be optimized with a new dimension and thus achieve enhanced performance compared to conventional systems with fixed antenna configuration.
Second, we derive a physical and EM based model for pixel antennas. Using microwave multiport network theory, we analyze the pixel antenna and model its configuration and characteristics such as radiation pattern as functions of a binary vector referred to as the antenna coder which represents the switch on and off states. Further, using the beamspace channel representation, we incorporate radiation patterns into the channel to derive a communication model for pixel antennas, where the channel can be controlled by antenna coder, allowing antenna coding optimization to enhance wireless systems.
Third, we optimize the antenna coding to maximize the channel gain in single-input single-output (SISO) pixel antenna systems. The antenna coding optimization is a binary optimization problem and can be solved by a successive exhaustive Boolean optimization approach. To reduce the computational complexity for optimization, we propose a codebook design for antenna coding and optimize the antenna coder by searching the codebook.
Fourth, we analyze the average channel gain of SISO pixel antenna systems and find that the upper bound of average channel gain is the effective aerial degrees-of-freedom of pixel antennas which can be found by singular value decomposition of the open-circuit radiation pattern matrix.
Fifth, we jointly optimize the antenna coding and transmit signal covariance matrix to maximize the channel capacity in MIMO pixel antenna systems by using exhaustive Boolean optimization approach and searching of the codebook design.
Sixth, we evaluate the performance of SISO and MIMO pixel antenna systems compared to conventional SISO and MIMO systems using antennas with fixed configuration. The simulation results show that using pixel antennas can enhance the average channel gain by up to 5.4 times and the channel capacity by up to 3.1 times, demonstrating the significant potential of pixel antennas in wireless communication systems.
Organization: Section II introduces the pixel antenna model and antenna coding. Section III presents the SISO pixel antenna system and antenna coding design for channel gain maximization. Section IV presents the MIMO pixel antenna system and antenna coding design for capacity maximization. Section V provides performance evaluation for the SISO and MIMO pixel antenna system. Section VI provides discussions and future work. Section VII concludes this work.
[…]
II Pixel Antenna
Figure 1: (a) Schematic for pixel antenna. (b) Multiport circuit network for pixel antenna. Red arrows represent the ports which replace switches.
Figure 2: Two examples of pixel antenna designs proposed in [8] and [10].
Figure 3: Model of the multiport circuit network for pixel antenna.
Figure 4: System diagram of the SISO pixel antenna system.
In this section, we introduce the pixel antenna model and the proposed antenna coding technique. As illustrated in the schematic in Fig. 1(a), the pixel antenna is based on discretizing a continuous radiation surface into small elements, which are referred to as pixels, and adjacent pixels can be connected through switches to flexibly control the antenna configuration and characteristics such as radiation pattern. Two examples of pattern-reconfigurable pixel antenna designs proposed in [8] and [10] are shown in Fig. 2.
The pixel antenna can be modeled using microwave multiport network theory [^1] The multiport circuit network model can be traced back to the method of moments (MoM) [31] which uses an impedance matrix to model an antenna. Using multiport circuit network to model lumped elements, such as capacitor, inductor, resistor, and RF switches, has been developed and demonstrated in [32], [33] and verified in reconfigurable antenna designs [7], [34].. Specifically, considering a pixel antenna which has $Q$ switches, we can model the pixel antenna as a multiport circuit network by replacing the $Q$ switches with $Q$ ports as shown in Fig. 1(b). Thus, the entire pixel antenna can be modeled as a $(Q+1)$-port circuit network, which consists of one antenna port and $Q$ pixel ports denoting the connections between adjacent pixels (which are numbered 2 to $Q+1$), as shown in Fig. 3. The $(Q+1)$-port circuit network can be characterized by a $(Q+1)\times(Q+1)$ impedance matrix $\mathbf{Z}$, which can be partitioned as
$\mathbf{Z}=\left[\begin{array}[]{ll}z_{\textrm{AA}}&\mathbf{z}_{\textrm{AP}}\\ \mathbf{z}_{\textrm{PA}}&\mathbf{Z}_{\textrm{PP}}\end{array}\right],$ (1)
where $z_{\textrm{AA}}\in\mathbb{C}$ is the self impedance for the antenna port, $\mathbf{Z}_{\textrm{PP}}\in\mathbb{C}^{Q\times Q}$ is the impedance matrix for the $Q$ pixel ports, $\mathbf{z}_{\textrm{AP}}\in\mathbb{C}^{1\times Q}$ is the trans-impedance relating the voltage of the antenna port with currents of the $Q$ pixel ports, and $\mathbf{z}_{\textrm{PA}}\in\mathbb{C}^{Q\times 1}$ is the transpose of $\mathbf{z}_{\textrm{AP}}$. We denote the voltages at the antenna port and pixel ports as $v_{\textrm{A}}\in\mathbb{C}$ and $\mathbf{v}_{\textrm{P}}=\left[v_{\textrm{P},1},\ldots,v_{\textrm{P},Q}\right]\in\mathbb{C}^{Q\times 1}$, respectively, and the currents at the antenna port and pixel ports as $i_{\textrm{A}}\in\mathbb{C}$ and $\mathbf{i}_{\textrm{P}}=\left[i_{\textrm{P},1},\ldots,i_{\textrm{P},Q}\right]\in\mathbb{C}^{Q\times 1}$, respectively. Using (1), we can relate the voltages and currents by
$\left[\begin{array}[]{c}v_{\textrm{A}}\\ \mathbf{v}_{\textrm{P}}\end{array}\right]=\left[\begin{array}[]{ll}z_{\textrm{AA}}&\mathbf{z}_{\textrm{AP}}\\ \mathbf{z}_{\textrm{PA}}&\mathbf{Z}_{\textrm{PP}}\end{array}\right]\left[\begin{array}[]{c}i_{\textrm{A}}\\ \mathbf{i}_{\textrm{P}}\end{array}\right].$ (2)
The switch connected to the $q$th pixel port has two states, on or off [^2] It is the RF switch that has on or off states and that it is not the pixel element that has on or off states. The RF switch in off state will only make the current at the RF switch zero, but this does not mean that the surface current through the pixel element will disappear, which is because pixel elements are closely coupled with each other. In addition, control of the RF switch states can be achieved by ensuring a dc path exists through the structure by using RF chokes.. As shown in Fig. 3, we can model the switch as a load impedance $z_{\textrm{L},q}$ $\forall q$ with open or short circuit, i.e. $z_{\textrm{L},q}=0$ or $\infty$. Therefore, a binary variable $b_{q}\in\left\{\textrm{0,1}\right\}$ can be used to denote the state of the switch connected to the $q$th port, i.e.
$z_{\textrm{L},q}=\begin{cases}0,&b_{q}=0,\>\textrm{i.e. switch\>on,}\\ \infty,&b_{q}=1,\>\textrm{i.e. switch\>off}.\end{cases}$ (3)
Numerically, for the switch off state, we can set $z_{\textrm{L},q}$ as a very large value, for example $z_{\textrm{L},q}=10^{8}$, to approach the infinite. We group $z_{\textrm{L},q}$ $\forall q$ into a diagonal load impedance matrix $\mathbf{Z}_{\textrm{L}}=\textrm{diag}(z_{\textrm{L},1},\ldots,z_{\textrm{L},Q})\in\mathbb{C}^{Q\times Q}$ and group $b_{q}$ $\forall q$ into a vector $\mathbf{b}=[b_{1},\ldots,b_{Q}]^{T}\in\mathbb{R}^{Q\times 1}$, which represents the states of all switches and is referred to as an antenna coder. Therefore, the load impedance matrix $\mathbf{Z}_{\textrm{L}}$ can be coded by $\mathbf{b}$, denoted as $\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)$, so that $\mathbf{v}_{\textrm{P}}$ and $\mathbf{i}_{\textrm{P}}$ can be related by
$\mathbf{v}_{\textrm{P}}=-\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)\mathbf{i}_{\textrm{P}}.$ (4)
Substituting (4) into (2), we can find that the currents at the pixel ports can be coded by the antenna coder $\mathbf{b}$, written as
$\mathbf{i}_{\textrm{P}}\left(\mathbf{b}\right)=-\left(\mathbf{Z}_{\textrm{PP}}+\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)\right)^{-1}\mathbf{z}_{\textrm{PA}}i_{\textrm{A}}.$ (5)
Accordingly, the radiation pattern of the pixel antenna can also be coded by the antenna coder $\mathbf{b}$. Given an antenna coder $\mathbf{b}$, we denote the radiation pattern of the pixel antenna excited by current $i_{\textrm{A}}$ as $\mathbf{e}\left(\mathbf{b}\right)=[\mathbf{e}_{\theta}^{T}\left(\mathbf{b}\right),\mathbf{e}_{\phi}^{T}\left(\mathbf{b}\right)]^{T}\in\mathbb{C}^{2K\times 1}$ where $\mathbf{e}_{\theta}\left(\mathbf{b}\right),\mathbf{e}_{\phi}\left(\mathbf{b}\right)\in\mathbb{C}^{K\times 1}$ are $\theta$ and $\phi$ polarization components over $K$ sampled spatial angles, respectively. The radiation pattern of the pixel antenna is the superposition of the radiation patterns of antenna and pixel ports weighted by currents. Thus we can express $\mathbf{e}\left(\mathbf{b}\right)$ as
$\mathbf{e}\left(\mathbf{b}\right)=\mathbf{e}_{\textrm{A}}i_{\textrm{A}}+\sum_{q=1}^{Q}\mathbf{e}_{\textrm{P},q}i_{\textrm{P},q}\left(\mathbf{b}\right)=\mathbf{E}_{\textrm{oc}}\mathbf{i}\left(\mathbf{b}\right),$ (6)
where $\mathbf{e}_{\textrm{A}}=[\mathbf{e}_{\textrm{A},\theta}^{T},\mathbf{e}_{\textrm{A},\phi}^{T}]^{T}\in\mathbb{C}^{2K\times 1}$ is the radiation pattern (with $\mathbf{e}_{\textrm{A},\theta},\mathbf{e}_{\textrm{A},\phi}\in\mathbb{C}^{K\times 1}$ being $\theta$ and $\phi$ polarization components) of the antenna port excited by a unit current when all the other ports are open-circuited, $\mathbf{e}_{\textrm{P},q}=[\mathbf{e}_{\textrm{P},q,\theta}^{T},\mathbf{e}_{\textrm{P},q,\phi}^{T}]^{T}\in\mathbb{C}^{2K\times 1}$ is the radiation pattern (with $\mathbf{e}_{\textrm{P},q,\theta},\mathbf{e}_{\textrm{P},q,\phi}\in\mathbb{C}^{K\times 1}$ being $\theta$ and $\phi$ polarization components) of the $q$th pixel port excited by unit current when all the other ports are open-circuit, the matrix $\mathbf{E}_{\textrm{oc}}=[\mathbf{e}_{\textrm{A}},\mathbf{e}_{\textrm{P},1},\ldots,\mathbf{e}_{\textrm{P},Q}]\in\mathbb{C}^{2K\times(Q+1)}$ collects the open-circuit radiation patterns of all ports and is named as open-circuit radiation pattern matrix, and $\mathbf{i}\left(\mathbf{b}\right)\in\mathbb{C}^{(Q+1)\times 1}$ collects the currents at all port, expressed as
$\mathbf{i}\left(\mathbf{b}\right)=\left[\begin{array}[]{c}i_{\textrm{A}}\\ \mathbf{i}_{\textrm{P}}\left(\mathbf{b}\right)\end{array}\right]=\left[\begin{array}[]{c}1\\ -\left(\mathbf{Z}_{\textrm{PP}}+\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)\right)^{-1}\mathbf{z}_{\textrm{PA}}\end{array}\right]i_{\textrm{A}}.$ (7)
The load impedance matrix $\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)$ can be coded by the antenna coder $\mathbf{b}$, which enables coding of the currents $\mathbf{i}\left(\mathbf{b}\right)$ and radiation pattern of the pixel antenna $\mathbf{e}\left(\mathbf{b}\right)$. Thus, given the impedance matrix $\mathbf{Z}$ and open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}$, we can find the radiation pattern of pixel antenna $\mathbf{e}\left(\mathbf{b}\right)$ configured by any antenna coder by (6) and (7) with low computational complexity of $\mathcal{O}\left(Q^{3}+Q^{2}+Q+2K(Q+1)\right)$. In Section V-D, we will show that this multiport circuit network model has the same accuracy as the full-wave EM simulation but requires significantly less computational time.
In total, there are $2^{Q}$ different combinations for the antenna coders and each antenna coder produces distinct radiation patterns. As a result, we can select the antenna coder $\mathbf{b}$ among a wide range of $2^{Q}$ combinations to find the optimum radiation pattern for wireless systems. In the following sections, we will introduce SISO and MIMO pixel antenna systems and show the antenna coding design for enhancing wireless systems.
III SISO Pixel Antenna System
As a starting point, we first introduce a SISO pixel antenna system, which is the simplest form for utilizing pixel antennas, to demonstrate the beamspace channel model, the antenna coding design, and codebook design for antenna coding.
III-A Beamspace Channel Model
The SISO pixel antenna system is equipped with a conventional antenna at the transmitter and a pixel antenna at the receiver, as shown in Fig. 4. To model the channel for the SISO pixel antenna system, we use the beamspace channel representation [35], [36]. Specifically, utilizing a virtual channel representation, we can represent the beamspace channel model in the angular domain, written as
$h\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}$ (8)
where $\mathbf{e}_{\textrm{T}}$ is the normalized radiation pattern of the conventional antenna at the transmitter satisfying $\left\|\mathbf{e}_{\textrm{T}}\right\|=1$, $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ is the normalized radiation pattern of the pixel antenna coded by antenna coder $\mathbf{b}_{\textrm{R}}$ at the receiver satisfying $\left\|\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)\right\|=1$, and $\mathbf{H}_{\textrm{V}}\in\mathbb{C}^{2K\times 2K}$ is the virtual channel matrix given by
$\mathbf{H}_{\textrm{V}}=\left[\begin{array}[]{cc}\mathbf{H}_{\textrm{V},\theta\theta}&\mathbf{H}_{\textrm{V},\theta\phi}\\ \mathbf{H}_{\textrm{V},\phi\theta}&\mathbf{H}_{\textrm{V},\phi\phi}\end{array}\right],$ (9)
where $\mathbf{H}_{\textrm{V},\theta\theta}$, $\mathbf{H}_{\textrm{V},\theta\phi}$, $\mathbf{H}_{\textrm{V},\phi\theta}$, $\mathbf{H}_{\textrm{V},\phi\phi}\in\mathbb{C}^{K\times K}$ are the virtual channel matrices for $\theta$ and $\phi$ polarizations, respectively, with each entry being the channel gain from an angle of departure (AoD) to an angle of arrival (AoA) among the $K$ spatial angles. We consider a rich scattering environment with Rayleigh fading and assume that $\left[\mathbf{H}_{\textrm{V}}\right]_{i,j}$ $\forall i,j$ are independent and identically distributed (i.i.d.) random variables following the complex Gaussian distribution $\mathcal{CN}\left(0,1\right)$.
From the beamspace channel model (8), we can see that the channel of SISO pixel antenna system $h\left(\mathbf{b}_{\textrm{R}}\right)$ can be coded by the antenna coder $\mathbf{b}_{\textrm{R}}$, which allows the antenna coding optimization to control the channel and enhance the system.
III-B Antenna Coding Design
For the SISO pixel antenna system, the received signal $y\in\mathbb{C}$ can be expressed as
$y=h\left(\mathbf{b}_{\textrm{R}}\right)x+n,$ (10)
where $x\in\mathbb{C}$ is the transmit signal and $n\in\mathbb{C}$ is additive white Gaussian noise following the complex Gaussian distribution $\mathcal{CN}\left(0,\sigma^{2}\right)$. Leveraging (8) and (10), the channel gain for the SISO pixel antenna system is given by
$\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}=\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (11)
Assuming the channel state information (CSI) for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ is perfectly known as the CSI can be estimated by using methods in [37], we aim to optimize the antenna coder for the pixel antenna $\mathbf{b}_{\textrm{R}}$ to maximize the channel gain, formulated as
$\underset{\mathbf{e}_{\textrm{R}},\mathbf{b}_{\textrm{R}},\mathbf{i}_{\textrm{R}},i_{\textrm{A}}^{\textrm{R}}}{\text{max}}$ $\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}$ (12)
s.t. $\left\|\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)\right\|=1,$ (13)
$\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{oc}}\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right),$ (14)
$\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\left[\begin{array}[]{c}1\\ -\left(\mathbf{Z}_{\textrm{PP}}+\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}_{\textrm{R}}\right)\right)^{-1}\mathbf{z}_{\textrm{PA}}\end{array}\right]i_{\textrm{A}}^{\textrm{R}},$ (17)
$\left[\mathbf{b}_{\textrm{R}}\right]_{q}\in\left\{0,1\right\},\forall q,$ (18)
where $\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ is the currents at all ports of the pixel antenna coded by $\mathbf{b}_{\textrm{R}}$ at the receiver and $i_{\textrm{A}}^{\textrm{R}}$ is the current at the antenna port. $i_{\textrm{A}}^{\textrm{R}}$ is an optimization variable to ensure that the constraint (13) can be satisfied for any given $\mathbf{b}_{\textrm{R}}$.
To handle the constraint (13), we normalize the radiation pattern $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ and therefore equivalently transform the problem (12)-(18) as
$\underset{\mathbf{e}_{\textrm{R}},\mathbf{b}_{\textrm{R}},\bar{\mathbf{i}}_{\textrm{R}}}{\text{max}\ \ }$ $\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}$ (19)
s.t. $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\frac{\mathbf{E}_{\textrm{oc}}\bar{\mathbf{i}}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)}{\left\|\mathbf{E}_{\textrm{oc}}\bar{\mathbf{i}}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)\right\|},$ (20)
$\bar{\mathbf{i}}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\left[\begin{array}[]{c}1\\ -\left(\mathbf{Z}_{\textrm{PP}}+\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}_{\textrm{R}}\right)\right)^{-1}\mathbf{z}_{\textrm{PA}}\end{array}\right],$ (23)
$\left[\mathbf{b}_{\textrm{R}}\right]_{q}\in\left\{0,1\right\},\forall q,$ (24)
where the variable $i_{\textrm{A}}^{\textrm{R}}$ is removed because it does not affect the normalized radiation pattern. Substituting (23) into (20), we can find the expression of $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ as a function of $\mathbf{b}_{\textrm{R}}$ and thus equivalently simplify the problem (19)-(24) as
$\underset{\mathbf{b}_{\textrm{R}}}{\text{max}\ \ }$ $\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}$ (25)
s.t. $\left[\mathbf{b}_{\textrm{R}}\right]_{q}\in\left\{0,1\right\},\forall q,$ (26)
which is an NP-hard binary optimization problem.
Various algorithms have been proposed and utilized for pixel antenna design and optimization such as the genetic algorithm [7], mixed integer linear programming [38], N-port characteristic mode analysis [39], [40], perturbation sensitivity analysis [41], the adjoint method combined with the method of moving asymptote [42], and reinforcement learning [43]. In this work, to solve the problem (25)-(26), we use an efficient algorithm called the Successive Exhaustive Boolean Optimization (SEBO) [44], which has two steps: 1) cyclically optimizing each block of the binary variables by exhaustive search until convergence, and 2) randomly flipping bits in the converged binary solution to check if there is any other local optimum with a better objective value. The computational complexity of SEBO is $\mathcal{O}\left(N_{\textrm{e}}2^{J}\right)$ where $J$ is the block size and $N_{\textrm{e}}$ is the number of iterations. Increasing $J$ can enhance the optimization performance. Particularly, when $J=Q$, the SEBO becomes the exhaustive search and global optimal solution can be found. However, increasing $J$ will exponentially increase the computational complexity. Thus, there is a trade-off between the computational complexity and optimization performance when using the SEBO to optimize antenna coding. More details on SEBO can be found in [44].
III-C Codebook Design for Antenna Coding
To reduce the computational complexity, we design a codebook for antenna coding optimization. Given a codebook which is defined as a set of $M$ different antenna coders $\mathcal{B}\triangleq\left\{\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{M}\right\}$, the channel gain is maximized by selecting the antenna coder from the codebook, formulated as
$\mathbf{b}_{\textrm{R}}^{\star}=\underset{\mathbf{b}_{\textrm{R}}\in\mathcal{B}}{\text{argmax}}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (27)
Given a codebook and channel realization of $\mathbf{H}_{\textrm{V}}$, we can optimize the antenna coder $\mathbf{b}_{\textrm{R}}$ by searching the codebook. However, it is important to design a good codebook to maximize ergodic channel gain, formulated as
$\underset{\mathbf{b}_{m}}{\text{max}}\ \$$\mathbb{E}\left[\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}\right]$ (28)
s.t. $\left[\mathbf{b}_{m}\right]_{q}\in\left\{0,1\right\},\forall m,q.$ (29)
The codebook design problem (28) and (29) is related to the vector quantization problem [45]. To handle the ergodic channel gain in (28), we leverage the sample average approximation method [46]. Specifically, we consider a training set containing $L$ realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ which is a random matrix with each entry following i.i.d. complex Gaussian distribution, given by
$\mathcal{H}\triangleq\left\{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mid l=1,\ldots,L\right\},$ (30)
where $\mathbf{H}_{\textrm{V}}^{\left[l\right]}$ denotes the $l$th virtual channel realization. Accordingly, we can approximate the expectation of channel gain in (28) as the sample average of channel gain, written as
$\mathbb{E}\left[\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}\right]\approx\frac{1}{L}\sum_{l=1}^{L}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (31)
Associated with each antenna coder, we can partition the training set $\mathcal{H}$ into $M$ subsets $\mathcal{H}_{1}$, $\mathcal{H}_{2}$, $\dots$, $\mathcal{H}_{M}$, where $\mathcal{H}_{m}$ is the neighborhood of the antenna coder $\mathbf{b}_{m}$ defined as
$\mathcal{H}_{m}$ $=\left\{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mid\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}\right$.
$\left.\qquad\hfill\geq\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m^{\prime}}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall l,\forall m^{\prime}\neq m\right\},\forall m.$ (32)
Making use of (31) and (32), we can equivalently reformulate the problem (28) and (29) as
$\underset{\mathbf{b}_{m}}{\text{max}}\ \$$\sum_{m=1}^{M}\sum_{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\in\mathcal{H}_{m}}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}$ (33)
s.t. $\left[\mathbf{b}_{m}\right]_{q}\in\left\{0,1\right\},\forall m,q.$ (34)
Note that to ensure that the sample average accurately approximates the ergodic channel gain, the entries of training set should be the independent realizations of random $\mathbf{H}_{\textrm{V}}$ and the size of training set $L$ should be large enough compared to the codebook size. Otherwise, such approximation is inaccurate and will degrade the codebook design.
We can see that the partition of training set $\mathcal{H}_{m}$ $\forall m$ relies on the antenna coder $\mathbf{b}_{m}$ $\forall m$ in the problem (33) and (34). Such coupled $\mathcal{H}_{m}$ and $\mathbf{b}_{m}$ make the problem (33) and (34) intractable. To address the problem (33) and (34), we follow the framework of the generalized Lloyd algorithm (GLA) [47] to optimize the training set partition $\mathcal{H}_{m}$ and antenna coder $\mathbf{b}_{m}$ alternatively via leveraging the nearest neighbor rule and the centroid condition. The methods are described next.
III-C1 Partition Optimization
One necessary condition for the optimal codebook is the nearest neighbor rule. Namely, all the channel realizations which have higher channel gain with the antenna coder $\mathbf{b}_{m}$ than with any other antenna coders should be assigned to the partition $\mathcal{H}_{m}$. Therefore, at iteration $i$ of the alternating optimization, we first optimize the partition by
$\mathcal{H}_{m}^{\left(i\right)}$ $=\left\{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mid\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}^{\left(i-1\right)}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}\right$.
$\left.\qquad\hfill\geq\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m^{\prime}}^{\left(i-1\right)}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall l,\forall m^{\prime}\neq m\right\},\forall m.$ (35)
where $\mathbf{b}_{m}^{\left(i-1\right)}$ is the antenna coder optimized at iteration i1i-1.
III-C2 Antenna Coder Optimization
The other necessary condition for the optimal codebook is the centroid condition. Namely, the optimal antenna coder $\mathbf{b}_{m}$ should be selected to maximize the average channel gain over the partition $\mathcal{H}_{m}$. Therefore, at iteration $i$, after the partition optimization, we then optimize the antenna coder by
$\mathbf{b}_{m}^{\left(i\right)}$ $=\underset{\left[\mathbf{b}_{m}\right]_{q}\in\left\{0,1\right\},\forall m,q}{\textrm{argmax}}\sum_{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\in\mathcal{H}_{m}^{\left(i\right)}}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall m,$ (36)
which is a binary optimization problem that can be solved by the SEBO algorithm.
By alternatively optimizing the partition (35) and the antenna coder (36), the average channel gain increases over each iteration so that the codebook design converges. Algorithm 1 summarizes the overall algorithm for codebook design.
Algorithm 1 Codebook Design for Antenna Coding
[…]
0: Codebook design $\mathcal{B}$;
1:Initialization: $i=0$, $\mathbf{b}_{m}^{\left(0\right)}$ $\forall m$;
2:repeat  $\text{\ \ \ \ \ \ }i=i+1;$
3: Partition Optimization:  $\text{\ \ \ \ \ \ }\text{Update}\ \mathcal{H}_{m}^{\left(i\right)}\ \forall m$ by (35) with $\mathbf{b}_{m}^{\left(i-1\right)}$;
4: Antenna Coder Optimization: $\text{\ \ \ \ \ \ }\text{Update}\ \mathbf{b}_{m}^{\left(i\right)}\ \forall m$ by solving (36) with $\mathcal{H}_{m}^{\left(i\right)}$;
5:until $\sum\left\|\mathbf{b}_{m}^{\left(i\right)}-\mathbf{b}_{m}^{\left(i-1\right)}\right\|/\sum\left\|\mathbf{b}_{m}^{\left(i\right)}\right\|\leq\epsilon$ or $i=i_{\textrm{max}}$;
6: Obtain $\mathbf{b}_{m}=\mathbf{b}_{m}^{\left(i\right)}$ $\forall m$;
7: Obtain $\mathcal{B}\triangleq\left\{\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{M}\right\}$
III-D Performance Analysis
Figure 5: System diagram of the MIMO pixel antenna system.
In order to obtain insights into the fundamental limits of pixel antennas, we derive the upper bound of the channel gain of SISO pixel antenna system as follows.
Using singular value decomposition (SVD), we can decompose the open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}$ as
$\mathbf{E}_{\textrm{oc}}=\mathbf{U}\mathbf{S}\mathbf{V}^{H}$ (37)
where $R=\textrm{rank}\left(\mathbf{E}_{\textrm{oc}}\right)$ is the rank of matrix $\mathbf{E}_{\textrm{oc}}$, $\mathbf{U}\in\mathbb{C}^{2K\times R}$ and $\mathbf{V}\in\mathbb{C}^{(Q+1)\times R}$ are semi-unitary matrices satisfying $\mathbf{U}^{H}\mathbf{U}=\mathbf{V}^{H}\mathbf{V}=\mathbf{I}$, and $\mathbf{S}=\textrm{diag}\left(s_{1},\ldots,s_{R}\right)\in\mathbb{R}^{R\times R}$ with $s_{i}$ representing the $i$th singular value. Making use of (6) and (37), we can rewrite $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ as
$\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{oc}}\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{U}\mathbf{S}\mathbf{V}^{H}\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right).$ (38)
This shows that any possible radiation pattern generated by the pixel antenna can be decomposed into $R$ orthogonal radiation patterns, i.e. the $R$ columns of $\mathbf{U}$. In the beamspace domain, the number of orthogonal radiation patterns that antennas can provide is defined as effective aerial degrees-of-freedom (EADoF) [48]. Hence, the EADoF of pixel antennas is given by $R$. Moreover, the singular value $s_{i}$ reflects the ability to radiate for the $i$th orthogonal radiation pattern. In other words, $s_{i}^{2}$ characterizes how much the power of the $i$th orthogonal radiation pattern is in the total radiated power. Thus, a high value of $s_{i}^{2}$ implies that the $i$th orthogonal radiation pattern is dominant in the total radiation pattern of the pixel antenna.
Accordingly, substituting (38) into (8), we can rewrite the beamspace channel of the SISO pixel antenna system as
$h\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{i}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{V}^{*}\mathbf{S}\mathbf{U}^{T}\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}=\mathbf{w}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\widetilde{\mathbf{h}}$, (39)
where $\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)\in\mathbb{C}^{R\times 1}$ is defined as
$\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{S}\mathbf{V}^{T}\mathbf{i}_{\textrm{R}}^{*}\left(\mathbf{b}_{\textrm{R}}\right),$ (40)
which satisfies $\left\|\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)\right\|=1$ due to the normalized radiation pattern $\left\|\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)\right\|=1$, and $\widetilde{\mathbf{h}}\in\mathbb{C}^{R\times 1}$ is defined as
$\widetilde{\mathbf{h}}=\mathbf{U}^{T}\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}},$ (41)
where $\widetilde{\mathbf{h}}]_{i}$ $\forall i$ are i.i.d. random variables following the complex Gaussian distribution $\mathcal{CN}\left(0,1\right)$ because of the $R$ orthogonal radiation patterns collected in $\mathbf{U}$.
From (39), we can observe that the channel of SISO pixel antenna system (with a conventional antenna at the transmitter and a pixel antenna at the receiver) is equivalent to a single-input multiple-output (SIMO) channel with combining. Therefore, we can find an upper bound for the channel gain of SISO pixel antenna system using Cauchy-Schwarz inequality
$\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}=\left|\mathbf{w}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\widetilde{\mathbf{h}}\right|^{2}\leq\left\|\widetilde{\mathbf{h}}\right\|^{2},$ (42)
where the upper bound is achieved if and only if the maximum ratio combiner is achieved, i.e.
$\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)=\frac{\widetilde{\mathbf{h}}}{\left\|\widetilde{\mathbf{h}}\right\|}$. (43)
However, it should be noted that for pixel antennas there is no guarantee that an antenna coder $\mathbf{b}_{\textrm{R}}$, such that $\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)$ satisfies (43), can be found. Therefore the upper bound for channel gain can be approached but cannot be guaranteed to be achieved. Accordingly, the upper bound for the average channel gain of SISO pixel antenna system is given by
$\mathbb{E}\left[\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}\right]\leq\mathbb{E}\left[\left\|\widetilde{\mathbf{h}}\right\|^{2}\right]=R.$ (44)
Therefore, this indicates that the average channel gain of the SISO pixel antenna system is upper bounded by how many orthogonal radiation patterns the pixel antenna can generate, i.e. the EADoF of pixel antennas, which is determined by the physical aperture of the pixel antenna [^3] It is difficult to write the relationship between EADoF and aperture in closed form as $\mathbf{E}_{\textrm{oc}}$ is obtained by simulating the pixel antenna by EM solver.. The average channel gain in comparison with the corresponding upper bound for SISO pixel antenna system with different physical apertures will be shown in Section V.
IV MIMO Pixel Antenna System
We introduce a MIMO pixel antenna system to demonstrate the beamspace channel model and antenna coding design for capacity maximization in this section.
IV-A Beamspace Channel Model
We consider a MIMO pixel antenna system. As shown in Fig. 5, the transmitter is equipped with $N_{\textrm{T}}$ pixel antennas, each of which is coded by $\mathbf{b}_{\textrm{T},n_{\textrm{T}}}\,\forall n_{\textrm{T}}$, and the receiver is equipped with $N_{\textrm{R}}$ pixel antennas, each of which is coded by $\mathbf{b}_{\textrm{R},n_{\textrm{R}}}\,\forall n_{\textrm{R}}$. We group $\mathbf{b}_{\textrm{T},n_{\textrm{T}}}\,\forall n_{\textrm{T}}$ into a matrix $\mathbf{B}_{\textrm{T}}=\left[\mathbf{b}_{\textrm{T},1},\ldots,\mathbf{b}_{\textrm{T},N_{\textrm{T}}}\right]\in\mathbb{R}^{Q\times N_{\textrm{T}}}$ and $\mathbf{b}_{\textrm{R},n_{\textrm{R}}}\,\forall n_{\textrm{R}}$ into a matrix $\mathbf{B}_{\textrm{R}}=\left[\mathbf{b}_{\textrm{R},1},\ldots,\mathbf{b}_{\textrm{R},N_{\textrm{R}}}\right]\in\mathbb{R}^{Q\times N_{\textrm{R}}}$. Utilizing the beamspace channel representation, the channel for MIMO pixel antenna system can be expressed as
$\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{R}}^{T}\left(\mathbf{B}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{E}_{\textrm{T}}\left(\mathbf{B}_{\textrm{T}}\right)$ (45)
where $\mathbf{E}_{\textrm{T}}\left(\mathbf{B}_{\textrm{T}}\right)\triangleq\text{[$\mathbf{e}_{\textrm{T},1}\left(\mathbf{b}_{\textrm{T},1}\right)$,$\ldots$,$\mathbf{e}_{\textrm{T},N_{\textrm{T}}}\left(\mathbf{b}_{\textrm{T},N_{\textrm{T}}}\right)$]}\in\mathbb{C}^{2K\times N_{\textrm{T}}}$ with $\mathbf{e}_{\textrm{T},n_{\textrm{T}}}\left(\mathbf{b}_{\textrm{T},n_{\textrm{T}}}\right)$ being the normalized radiation pattern of the $n_{\textrm{T}}$th transmit pixel antenna satisfying $\left\|\mathbf{e}_{\textrm{T},n_{\textrm{T}}}\left(\mathbf{b}_{\textrm{T},n_{\textrm{T}}}\right)\right\|=1$, $\mathbf{E}_{\textrm{R}}\left(\mathbf{B}_{\textrm{R}}\right)\triangleq[\mathbf{e}_{\textrm{R},1}\left(\mathbf{b}_{\textrm{R},1}\right),\ldots,\mathbf{e}_{\textrm{R},N_{\textrm{R}}}\left(\mathbf{b}_{\textrm{R},N_{\textrm{R}}}\right)]\in\mathbb{C}^{2K\times N_{\textrm{R}}}$ with $\mathbf{e}_{\textrm{R},n_{\textrm{R}}}\left(\mathbf{b}_{\textrm{R},n_{\textrm{R}}}\right)$ being the normalized radiation pattern of the $n_{\textrm{R}}$th receive pixel antenna satisfying $\left\|\mathbf{e}_{\textrm{R},n_{\textrm{R}}}\left(\mathbf{b}_{\textrm{R},n_{\textrm{R}}}\right)\right\|=1$, and $\mathbf{H}_{\textrm{V}}\in\mathbb{C}^{2K\times 2K}$ is the virtual channel matrix given by (9). Similarly, from the beamspace channel model (45), we can see that the channel of MIMO pixel antenna system $\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)$ can be coded by the transmit and receive antenna coders $\mathbf{B}_{\textrm{T}}$ and $\mathbf{B}_{\textrm{R}}$, which allows antenna coding optimization to enhance the MIMO system.
IV-B Channel Capacity
For the MIMO pixel antenna system, the received signal $\mathbf{y}\in\mathbb{C}^{N_{\textrm{R}}\times 1}$ can be expressed as
$\mathbf{y}=\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{x}+\mathbf{n},$ (46)
where $\mathbf{x}\in\mathbb{C}^{N_{\textrm{T}}\times 1}$ is the transmit signal with covariance matrix given by $\mathbf{X}=\mathbb{E}\left[\mathbf{x}\mathbf{x}^{H}\right]$ and $\mathbf{n}\in\mathbb{C}^{N_{\textrm{R}}\times 1}$ is the additive white Gaussian noise following the complex Gaussian distribution $\mathcal{CN}\left(\boldsymbol{0},\sigma^{2}\mathbf{I}\right)$. Accordingly, the channel capacity for the MIMO pixel antenna system is given by
$C\left(\mathbf{X},\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)=\log_{2}\left|\mathbf{I}+\frac{1}{\sigma^{2}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{X}\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|.$ (47)
Assuming the CSI for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ is perfectly known, we aim to jointly optimize the transmit signal covariance matrix $\mathbf{X}$ and the transmit and receive antenna coders $\mathbf{B}_{\textrm{T}}$ and $\mathbf{B}_{\textrm{R}}$ to maximize the channel capacity for the MIMO pixel antenna system, which can be formulated as
$\underset{\mathbf{X},\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}$ $\log_{2}\left|\mathbf{I}+\frac{1}{\sigma^{2}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{X}\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|$ (48)
s.t. $\textrm{Tr}\left(\mathbf{X}\right)\leq P,$ (49)
$\left[\mathbf{B}_{\textrm{T}}\right]_{i,j}\in\left\{0,1\right\},\forall i,j,$ (50)
$\left[\mathbf{B}_{\textrm{R}}\right]_{i,j}\in\left\{0,1\right\},\forall i,j,$ (51)
where $P$ denotes the maximum transmit power.
IV-C Antenna Coding Design
To solve the capacity maximization problem (48)-(51), in the following, we consider two cases of the transmit signal covariance matrix.
IV-C1 Antenna Coding Design with Uniform Power Allocation
We first consider when the uniform power allocation is utilized in the MIMO pixel antenna system, i.e. $\mathbf{X}=\frac{P}{N_{\textrm{T}}}\mathbf{I}$. In this case, we only need to optimize the transmit and receive antenna coders $\mathbf{B}_{\textrm{T}}$ and $\mathbf{B}_{\textrm{R}}$ for capacity maximization, formulated as
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\log_{2}\left|\mathbf{I}+\frac{P}{\sigma^{2}N_{\textrm{T}}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|$ (52)
s.t. $\left[\mathbf{B}_{\textrm{T}}\right]_{i,j}\in\left\{0,1\right\},\>\forall i,j,$ (53)
$\left[\mathbf{B}_{\textrm{R}}\right]_{i,j}\in\left\{0,1\right\},\>\forall i,j,$ (54)
which is an NP-hard binary optimization problem. Similar to the channel gain maximization in the SISO pixel antenna, we use the SEBO algorithm to solve the problem (52)-(54).
IV-C2 Joint Antenna Coding and Waterfilling Design
We next consider joint antenna coding and transmit signal covariance matrix design for capacity maximization in the MIMO pixel antenna system. For the channel matrix $\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)$ with any given transmit and receive antenna coders $\mathbf{B}_{\textrm{T}}$ and $\mathbf{B}_{\textrm{R}}$, it is well known that the transmit signal covariance matrix design based on waterfilling power allocation provides the maximum channel capacity [49], i.e.
$\sum_{i=1}^{N_{\textrm{min}}}\log_{2}\left(1+\frac{P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}{\sigma^{2}}\right),$ (55)
where $N_{\textrm{min}}=\textrm{min}\left(N_{\textrm{T}},N_{\textrm{R}}\right)$, $\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)$ is the $i$th eigenvalue of the matrix $\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)$, and $P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)$ represents the waterfilling power allocation expressed as
$P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)=\left(\mu-\frac{\sigma^{2}}{\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}\right)^{+},$ (56)
with $\mu$ chosen to satisfy the total transmit power constraint
$\sum_{i=1}^{N_{\textrm{min}}}P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)=P.$ (57)
As a result, the channel capacity with waterfilling power allocation is also a function of $\mathbf{B}_{\textrm{T}}$ and $\mathbf{B}_{\textrm{R}}$, so that the antenna coding design can be formulated as
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\sum_{i=1}^{N_{\textrm{min}}}\log_{2}\left(1+\frac{P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}{\sigma^{2}}\right)$ (58)
s.t. $\left[\mathbf{B}_{\textrm{T}}\right]_{i,j}\in\left\{0,1\right\},\>\forall i,j,$ (59)
$\left[\mathbf{B}_{\textrm{R}}\right]_{i,j}\in\left\{0,1\right\},\>\forall i,j,$ (60)
which is an NP-hard binary optimization problem. Similarly, we use the SEBO algorithm to solve the problem (58)-(60).
IV-D Antenna Coding Design with Codebook
To reduce computational complexity, we use the codebook proposed in Section III-C to facilitate antenna coding design in the MIMO pixel antenna system. Specifically, given the codebook $\mathcal{B}\triangleq\left\{\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{M}\right\}$, we select the antenna coder for each transmit and receive pixel antenna from the codebook, i.e. $\mathbf{b}_{\textrm{T},n_{\textrm{T}}},\mathbf{b}_{\textrm{R},n_{\textrm{R}}}\in\mathcal{B}$ $\forall n_{\textrm{T}},n_{\textrm{R}}$, to maximize the channel capacity of MIMO pixel antenna system. For the uniform power allocation case, we formulate the channel capacity maximization as
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\log_{2}\left|\mathbf{I}+\frac{P}{\sigma^{2}N_{\textrm{T}}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|$ (61)
s.t. $\left[\mathbf{B}_{\textrm{T}}\right]_{:,n_{\textrm{T}}}\in\mathcal{B},\>\forall n_{\textrm{T}},$ (62)
$\left[\mathbf{B}_{\textrm{R}}\right]_{:,n_{\textrm{R}}}\in\mathcal{B},\>\forall n_{\textrm{R}}.$ (63)
For the waterfilling power allocation case, we formulate the channel capacity maximization as
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\sum_{i=1}^{N_{\textrm{min}}}\log_{2}\left(1+\frac{P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}{\sigma^{2}}\right)$ (64)
s.t. $\left[\mathbf{B}_{\textrm{T}}\right]_{:,n_{\textrm{T}}}\in\mathcal{B},\>\forall n_{\textrm{T}},$ (65)
$\left[\mathbf{B}_{\textrm{R}}\right]_{:,n_{\textrm{R}}}\in\mathcal{B},\>\forall n_{\textrm{R}}.$ (66)
Both the problem (61)-(63) and problem (64)-(66) can be solved by successively searching the codebook for each transmit and receive pixel antenna.
V Performance Evaluation
We evaluate the performance of SISO pixel antenna system and MIMO pixel antenna system with the proposed antenna coding design in this section.
V-A Simulation Setup and Pixel Antenna Design
We consider a propagation environment with rich scattering, which has a 2-D uniform power angular spectrum [^4] Herein we consider 2-D uniform power angular spectrum for illustration. It should be noted that other power angular spectrum also applies., i.e. uniform over the azimuth angle on the XOY plane, with equally likely polarization. We set the angular resolution as $\Delta\phi=5\text{\textdegree}$ and thus we have $K=72$ sampled angles. For the SISO pixel antenna system, we assume that the transmit antenna is fixed with an ideal isotropic radiation pattern and the receive antenna uses a pixel antenna. For the MIMO pixel antenna system, both transmit and receive antennas use pixel antennas.
We consider two examples of pixel antennas operating at 2.4 GHz where the wavelength is $\lambda=125$ mm. The two pixel antennas are designed based on discretizing the radiating patch surface of conventional microstrip patch antennas [^5] Herein we consider the pixelated microstrip patch antenna for illustration. However, it should be noted that pixel antenna is general and can be designed based on discretizing the continuous radiating surface of different types of antennas such as monopole antenna, not restricted to microstrip patch antenna. into a grid of pixels, but having different physical apertures, $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$. The geometry and dimension for the two pixel antennas are illustrated in Fig. 6 and summarized in Table I. For both pixel antennas, there are in total 40 ports, including 1 antenna port and $Q=39$ pixel ports, and $\times5$ pixels with pixel size being $5\textrm{mm}\times 5\textrm{mm}$ and $11\textrm{mm}\times 11\textrm{mm}$, respectively. The pixelized patch and ground plane are printed on the Rogers 4003C substrate, where the thickness is 1.524 mm, the permittivity is 3.55, and the loss tangent is 0.0027. We use CST studio suite, a full-wave EM solver, to simulate the $(Q+1)$-port circuit network of pixel antennas to achieve the impedance matrix $\mathbf{Z}\in\mathbb{C}^{(Q+1)\times(Q+1)}$ as well as the open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}\in\mathbb{C}^{2K\times(Q+1)}$.
Figure 6: Pixel antennas with different physical apertures (a) $0.25\lambda\times 0.25\lambda$ and (b) $0.5\lambda\times 0.5\lambda$. Red arrows represent the antenna ports and blue arrows represent the pixel ports.
TABLE I: Geometry and Dimensions for Pixel Antennas
Physical aperture $0.25\lambda\times 0.25\lambda$ $0.5\lambda\times 0.5\lambda$
Size of ground $\times31.25mm$ $\times62.5mm$
Size of patch $\times29mm$ $\times59mm$
Number of pixels $\times5$ $\times5$
Size of pixel $\times5mm$ $\times11mm$
Gap between pixels 1mm 1mm
Patch and ground spacing 12.5mm 12.5mm
Figure 7: Singular values $s_{i}$ $\forall i$ of open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}$ and cumulative distribution function of $s_{i}^{2}$ for pixel antennas with different physical apertures (a) $0.25\lambda\times 0.25\lambda$ and (b) $0.5\lambda\times 0.5\lambda$.
The singular values $s_{i}$ $\forall i$ (sorted in descending order) of the open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}$ for pixel antennas with different physical apertures are shown in Fig. 7. From Fig. 7, we can observe that the majority of the singular values are almost zero. For example, more than half of the singular values are smaller than 0.1. Recall that the singular value $s_{i}$ reflects the ability to radiate for the $i$th orthogonal radiation pattern, i.e. $s_{i}^{2}$ characterizes how much the power of the $i$th orthogonal radiation pattern is in the total radiated power. Thus, a small value of $s_{i}^{2}$ implies that the $i$th orthogonal radiation pattern makes negligible contribution to the total radiation pattern of the pixel antennas. As a result, to identify the EADoF of the pixel antennas $R=\textrm{rank}\left(\mathbf{E}_{\textrm{oc}}\right)$, i.e. how many orthogonal radiation patterns the pixel antennas can generate, we consider the cumulative distribution function of $s_{i}^{2}$, defined as
$F_{i}=\frac{\sum_{j=1}^{i}s_{j}^{2}}{\sum_{j=1}^{Q+1}s_{j}^{2}},$ (67)
which characterizes how much power the first $i$ orthogonal radiation patterns contribute to the total radiated power. The cumulative distribution function of $s_{i}^{2}$ for pixel antennas with different physical apertures are also shown in Fig. 7. We can see that the cumulative distribution function increases with the singular value index and approaches unity after a certain index. Based on this, we select the EADoF of pixel antennas as
$R=\underset{F_{i}\geq T}{\textrm{argmin}}\>i,$ (68)
which implies that the contribution of $R$ orthogonal radiation patterns to the total radiated power is more than the threshold $T$. In this work, we set the threshold $T$ as 0.998, so that the EADoF of the pixel antennas with the physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are 5 and 7, respectively. This shows that the EADoF of pixel antennas increases with the physical aperture. Moreover, recall that the upper bound for the average channel gain of SISO pixel antenna system is given by the EADoF of pixel antennas. The average channel gain of SISO pixel antenna system in comparison with its upper bound will be demonstrated in the following subsection.
V-B SISO Pixel Antenna System Performance
We first evaluate the performance of SISO pixel antenna system. Utilizing Monte Carlo method, we generate multiple realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ and for each realization, we use the SEBO algorithm [44], with block size of $J=10$ (which achieves a good optimization performance with an acceptable complexity), to solve the problem (25)-(26) so as to maximize the channel gain $\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}$ with perfect CSI. In addition, we also maximize the channel gain with the codebook design for antenna coding introduced in Section III-C. The number of virtual channel realizations in the training set is $L=15000$. The antenna coders in the codebook design with size [^6] Due to limited space, it would be messy to show the antenna coders in the codebook with size larger than 8, so we omit them. of 2, 4, and 8 are listed in Table II. The number of realizations in Monte Carlo method is 100, which is enough to obtain an accurate average performance. The average channel gain of the SISO pixel antenna system optimized with perfect CSI and with the codebook is shown in Fig. 8. The two pixel antennas with different physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are considered and the conventional SISO system is also compared as the benchmark. From Fig. 8, we can make three observations.
TABLE II: Antenna Coders in the Codebook with Size of 2, 4, 8
Physical aperture $0.25\lambda\times 0.25\lambda$ Physical aperture $0.5\lambda\times 0.5\lambda$
Codebook 000110100000011101101011000001101001101 010000110010110101101011001010011001000
size = 2 101010111110110111010001001111101010101 100010101010010110110011100110000001101
101001011100010111110010000110110001000 001010110001100101101100100100000101010
Codebook 110111011110010100110110111000101010110 111000011111110101000001000110100101111
size = 4 110100000010001000110111110000011011110 100101111010010100101000011010001101110
010110001001000101111011010111111101110 100101000010001001000010000101111100101
000000111110010010110111101001011101101 111111010000001000101111110100110110000
000111000011011101111000110111010001011 000001101000100110101010101100001000011
010011110101110101100101111011100100001 111111000111000100011000100000110001011
Codebook 001011100000010001110110100110110101100 010000110101010111010100100001001100000
size = 8 010111010110011100011011111011111101010 100100100010000000010110000101111000011
000100101101010000110110100110000101001 010111110100011100001111100100010100111
001010011111100011010111101101100000000 000101001111111101010101110100111100011
100001010010000011111000110010010111110 110000011011011000100001010010100001001
Figure 8: Average channel gain of the SISO pixel antenna system with physical apertures (a) $0.25\lambda\times 0.25\lambda$ and (b) $0.5\lambda\times 0.5\lambda$, optimized with perfect CSI and with codebook.
Figure 9: Channel capacity of the conventional MIMO system and the MIMO pixel antenna system with the uniform power allocation (UP) and waterfilling (WF) optimized with perfect CSI.
Firstly, compared with the conventional SISO system which has a unit average channel gain from the complex Gaussian distribution $\mathcal{CN}\left(0,1\right)$, the SISO pixel antenna system can significantly enhance the channel gain, demonstrating the benefits of pixel antennas. Specifically, the average channel gain with physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ optimized with the perfect CSI are 4 and 5.4, respectively. The channel gain enhanced by SISO pixel antenna system can be explained by noting that the pixel antenna can flexibly adjust its configuration so that its radiation pattern can adapt to the virtual channel matrix to coherently add the multiple paths in the propagation environment.
Secondly, the average channel gain of SISO pixel antenna system optimized with the codebook increases with the codebook size. This is because increasing the codebook size can generate more antenna coders, so that the pixel antenna becomes more flexible and provides more diverse radiation patterns to enhance the channel. When the codebook size is large enough, the average channel gain optimized with the codebook provides performance similar to that optimized with perfect CSI. However the computational complexity for optimization can be significantly reduced. Particularly, even when the codebook size is only 2, utilizing the codebook design can still enhance the average channel gain by around 50% for different physical apertures. Therefore, this demonstrates the benefit of codebook design for pixel antennas.
Thirdly, for the SISO pixel antenna system, increasing the physical aperture is beneficial to enhance channel gain. For example, when optimized with perfect CSI, the channel gain with physical aperture $0.5\lambda\times 0.5\lambda$ is 35% higher than that with $0.25\lambda\times 0.25\lambda$. This is because pixel antennas with larger physical aperture can provide more orthogonal radiation patterns, i.e. larger EADoF. Specifically, the EADoF for the pixel antenna with physical aperture $0.25\lambda\times 0.25\lambda$ is 5 while the EADoF with $0.5\lambda\times 0.5\lambda$ is 7. Therefore, pixel antennas with larger physical aperture achieve a higher upper bound for the average channel gain (44), resulting in a higher average channel gain. Moreover, it is worth noting that the upper bound of the average channel gain of SISO pixel antenna system (44) is generally hard to achieve. For example, for the physical aperture $0.25\lambda\times 0.25\lambda$, the average channel gain optimized with perfect CSI is 4 while the upper bound is 5. This is because it is not guaranteed to find the antenna coder $\mathbf{b}_{\textrm{R}}$ such that the pixel antenna current $\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ and the $\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)$ given by (40) satisfy the condition of maximum ratio combiner (43).
Overall, we have shown that using pixel antennas is beneficial to enhance the channel gain of SISO system.
V-C MIMO Pixel Antenna System Performance
Figure 10: Channel capacity of the MIMO pixel antenna system optimized with codebook.
We next evaluate the performance of MIMO pixel antenna systems. Similarly, we utilize the Monte Carlo method to generate multiple realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$. For each realization, we optimize the antenna coding design with uniform power allocation by solving the problem (52)-(54) and also jointly optimize the antenna coding and waterfilling design by solving the problem (58)-(60), for maximizing the channel capacity of MIMO pixel antenna system with perfect CSI. The SEBO algorithm [44], with block size $J=10$, is utilized to solve the two problems. The channel capacity of the MIMO pixel antenna system with uniform power allocation (UP) and waterfilling (WF) optimized with perfect CSI is shown in Fig. 9. The channel capacity is averaged over multiple channel realizations. The two pixel antennas with different physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are considered and the conventional MIMO system with UP and WF are compared. Different numbers of transmit and receive antennas $N_{\textrm{T}}$ and $N_{\textrm{R}}$ are also considered for the MIMO pixel antenna system and conventional MIMO system. From Fig. 9, we can make four observations.
Firstly, in comparison with the conventional MIMO system, the MIMO pixel antenna system can effectively enhance channel capacity, demonstrating the benefits of pixel antennas. For example, in the case of $N_{\textrm{T}}=2$ and $N_{\textrm{R}}=2$, the channel capacity with UP at low SNR of 0 dB for the conventional MIMO system and the MIMO pixel antenna system with physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ is 1.75, 4.50, and 5.38 bps/Hz, respectively, which implies that the capacity for MIMO pixel antenna systems are 2.6 and 3.1 times higher than that for the conventional MIMO system. On the other hand, in the same case but at high SNR of 30 dB, the channel capacity with UP for the conventional MIMO system and the MIMO pixel antenna system with physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are 17.48, 23.58 and 24.65 bps/Hz, respectively, which means that using pixel antenna can enhance the capacity by 35% and 41%. The channel capacity enhanced by MIMO pixel antenna system occurs because the multiple pixel antennas at the transmitter and receivers can flexibly adjust their radiation patterns to adapt to the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ so that the multiple paths with different AoA and AoD in the propagation environment can be coherently added to enhance the channel $\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)$ in (45) as well as the capacity.
Secondly, increasing the physical aperture is beneficial to enhance the channel capacity of MIMO pixel antenna system. For example, at low SNR of 0 dB, the channel capacity with UP for MIMO pixel antenna system with physical apertures $0.5\lambda\times 0.5\lambda$ is 20% higher than that with $0.25\lambda\times 0.25\lambda$. This is because pixel antennas with larger physical aperture provide larger EADoF which increases the channel gain between each pair of transmit and receive pixel antennas and therefore leads to higher channel capacity.
Thirdly, for MIMO pixel antenna system, increasing the number of transmit antennas and receive antennas is beneficial to improve the channel capacity. For example, with the same physical aperture, the channel capacity at high SNR of 30 dB for the $4\times 4$ MIMO pixel antenna system is nearly twice that for the $2\times 2$ MIMO pixel antenna system, which is consistent with the conventional MIMO system. That is to say, utilizing pixel antennas to construct MIMO systems will not affect the multiplexing gain of MIMO system. Moreover, it will bring extra channel gain on top of the multiplexing gain.
Fourthly, the joint antenna coding and waterfilling design achieves higher channel capacity than antenna coding design with uniform power allocation, especially when there are more transmit antennas than receive antennas. For example, for the $4\times 2$ MIMO pixel antenna system with physical aperture $0.25\lambda\times 0.25\lambda$, the channel capacity with UP and WF is 4.39 and 5.99 bps/Hz at low SNR of 0 dB, respectively. This implies that joint antenna coding and waterfilling design can increase the capacity by 36%. Thus, this shows that pixel antennas can synergize with waterfilling design to enhance channel capacity in MIMO systems.
We also use the codebook to optimize the antenna coding design with uniform power allocation and waterfilling to maximize the channel capacity of MIMO pixel antenna system. Specifically, we solve the problem (61)-(63) and problem (64)-(66) by successively searching the codebook for each transmit and receive pixel antenna. The channel capacity of the MIMO pixel antenna system with UP and WF optimized with the codebook is shown in Fig. 10. The channel capacity is averaged over channel realizations and different physical apertures are considered (we only show the $4\times 2$ case for brevity). From Fig. 10, we can make two observations.
Firstly, the channel capacity of MIMO pixel antenna system optimized with the codebook increases with the physical aperture. Joint antenna coding and waterfilling design achieves higher channel capacity than antenna coding design with uniform power allocation. These are consistent with the case optimized with CSI.
Secondly, the channel capacity of MIMO pixel antenna system optimized with the codebook increases with the codebook size, which is because increasing the codebook size gives the pixel antenna more diverse radiation patterns to enhance the channel. When the codebook size is large enough, the capacity optimized with codebook is similar to that optimized with perfect CSI but the computational complexity can be reduced. Even when the codebook size is only 2, using the codebook design is still beneficial. For example, the capacity with UP for the $4\times 2$ MIMO pixel antenna system at low SNR of 0 dB is 30% higher than that for conventional MIMO system when codebook size is 2. Thus, this shows the benefit and effectiveness of codebook for MIMO pixel antenna system.
Overall, we have shown that using pixel antennas is beneficial to enhance the channel capacity of MIMO system.
V-D Evaluation of Multiport Circuit Network Model
To validate the accuracy of the multiport circuit network model, we compare it with the full-wave EM simulation using CST studio suite. Specifically, in CST studio suite, we simulate the pixel antenna with physical aperture $0.5\lambda\times 0.5\lambda$ configured by different antenna coders in the codebook (with codebook size of 4). The radiation patterns of the pixel antenna, obtained by the EM simulation and the multiport circuit network model (6) and (7), are shown in Fig. 11. We can find that the EM simulated radiation patterns are almost same as those computed by the multiport circuit network model for different antenna coders, showing the accuracy of the proposed model. In addition, the radiation patterns for different antenna coders are diverse, covering different angle ranges, which is consistent with the fact that the codebook is designed to maximize the ergodic channel gain in a multipath propagation environment with 2-D uniform power angular spectrum.
More importantly, under the same accuracy, the multiport circuit network model is significantly more computationally efficient. We measure the computational time to calculate the radiation pattern using the EM simulation (CST studio suite) and the proposed multiport circuit network model, respectively, in a laptop (with AMD Ryzen 9 7940HX CPU and 16 GB memory). The average computational time for the EM simulation and the multiport circuit network model, denoted as $T_{\textrm{EM}}$ and $T_{\textrm{circuit}}$, is measured around 56.59s and 0.000535s, respectively, which shows that the proposed model is around $10^{5}$ time faster than the EM simulation.
[…]
Last but not least, we completely compare the computational time for the whole process to optimize the antenna coder using the proposed multiport circuit network model and the EM simulation in the following two cases.
First, when using SEBO to optimize the antenna coder, i.e. solving the problem (25)-(26), the total number of calculations of radiation pattern is $N_{\textrm{e}}2^{J}$, so the total computational time for the proposed multiport circuit network model is
$T_{\textrm{total,circuit}}=T_{0}+N_{\textrm{e}}2^{J}T_{\textrm{circuit}},$ (69)
while the total computational time using EM simulation is
$T_{\textrm{total,EM}}=N_{\textrm{e}}2^{J}T_{\textrm{EM}}.$ (70)
If considering $N_{\textrm{e}}=4$ (the least number to optimize all the $Q$ switches), we have $T_{\textrm{total,circuit}}=\textrm{722}\,\textrm{s}$ and $T_{\textrm{total,EM}}=\textrm{231792}\,\textrm{s}.$
Second, when using the codebook to optimize the antenna coder, we need to design the codebook, which needs $N_{\textrm{c}}=N_{\textrm{o}}\left(\sum_{m=1}^{M}N_{\textrm{e},m}2^{J}\right)$ calculations of radiation pattern, where $N_{\textrm{o}}$ is the number of iterations of antenna coder optimization in GLA and $N_{\textrm{e},m}$ is the number of iterations for using SEBO to optimize the $m$th antenna coder in codebook with size of $M$. Once the codebook is ready, we search the codebook by $M$ calculations of radiation pattern. Totally, we need $N_{\textrm{c}}+M$ calculations of radiation pattern to use codebook to optimize antenna coder, so the total computational time for the proposed multiport circuit network model is
$T_{\textrm{total,circuit}}=T_{0}+\left(N_{\textrm{o}}\left(\sum_{m=1}^{M}N_{\textrm{e},m}2^{J}\right)+M\right)T_{\textrm{circuit}},$ (71)
while the total computational time using EM simulation is
$T_{\textrm{total,EM}}=\left(N_{\textrm{o}}\left(\sum_{m=1}^{M}N_{\textrm{e},m}2^{J}\right)+M\right)T_{\textrm{EM}}.$ (72)
If considering $N_{\textrm{o}}=1$ (the least number to perform antenna coding optimization in GLA), $N_{\textrm{e},m}=4$ $\forall m$, and e.g. $M=16$, we have $T_{\textrm{total,circuit}}=\textrm{755}\,\textrm{s}$ and $T_{\textrm{total,EM}}=\textrm{3709587}\,\textrm{s}$.
From the above complete comparison, we have shown that the proposed multiport circuit network model is significantly more computationally efficient than the EM simulation.
Figure 11: Radiation patterns of the pixel antenna (with physical aperture $0.5\lambda\times 0.5\lambda$) configured by antenna coders in the codebook (with codebook size of 4) using EM simulation and multiport circuit network model. Unit: dB.
VI Discussions and Future Work
VI-A Continuous Antenna Coding
It is possible to continuously adjust the load impedance $z_{\textrm{L},q}$ to implement a continuous antenna coding. To that end, we can use variable reactive load, such as varactor, to replace the RF switch in the pixel antenna, so that $z_{\textrm{L},q}$ can be modeled as $z_{\textrm{L},q}=jx_{\textrm{L},q}$, where $j$ is the imaginary unit and $x_{\textrm{L},q}$ is the continuous variable reactance. Accordingly, the continuous antenna coding design becomes a continuous optimization problem, which is easier to solve and can provide better performance compared with the binary antenna coding. However, implementing the continuous antenna coding increases the circuit complexity to control the pixel antenna since continuous dc bias voltages are required. Therefore, there is a tradeoff between the circuit complexity, system performance, and difficulty for optimization.
VI-B Antenna Coding Optimization for More Scenarios
Antenna coding can be optimized for more scenarios in future research, including, but are not limited to, the following:
VI-B1 Multi-User MIMO Pixel Antenna System
We can jointly optimize the precoding, combining, and antenna coding at transmitter and users to enhance the signal-to-interference-plus-noise ratio to maximize the weighted sum rate.
VI-B2 OFDM MIMO Pixel Antenna System
We can derive the frequency dependent model for pixel antenna and subsequently jointly optimize the precoding at each carrier frequency and the antenna coding of pixel antenna to maximize the capacity.
VI-B3 Pixel Antenna System for Wireless Power Transfer
We can jointly optimize the waveform, beamforming, and antenna coding to maximize the output dc power.
VI-C Pixel Antenna Element Spacing
There is a limitation on the spacing between adjacent pixel antenna elements, which should be neither too small nor too large, as explained in the following.
VI-C1 Inter-Element Electromagnetic Interference
When the spacing is too small, mutual coupling between the pixel antennas will be high and degrade the performance. In this work, we address the inter-element electromagnetic interference (EMI) issue in the MIMO pixel antenna system through separating the pixel antennas by a suitable distance to suppress the mutual coupling. For example, we separate two pixel antennas with physical aperture $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ by half-wavelength and one wavelength, respectively. Moreover, we can extend the proposed multiport circuit network model to take the inter-element EMI issue into account. For example, we consider two closely placed pixel antennas, which can be modeled as a ($2Q+2$)-port circuit network characterized by a $(2Q+2)\times(2Q+2)$ impedance matrix $\mathbf{Z}$. Given the antenna coders for pixel antennas 1 and 2, denoted as $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$, we can find the corresponding load impedance $\mathbf{Z}_{\textrm{L},1}\left(\mathbf{b}_{1}\right)$ and $\mathbf{Z}_{\textrm{L},2}\left(\mathbf{b}_{2}\right)$. Thus, from $\mathbf{Z}$, $\mathbf{Z}_{\textrm{L},1}\left(\mathbf{b}_{1}\right)$, and $\mathbf{Z}_{\textrm{L},2}\left(\mathbf{b}_{2}\right)$, we can find the currents $\mathbf{i}_{1}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ and $\mathbf{i}_{2}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ and radiation patterns $\mathbf{e}_{1}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ and $\mathbf{e}_{2}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ for pixel antennas 1 and 2. Using the beamspace channel model, we can link the radiation patterns to the channel, so that we can optimize $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ to enhance the performance. By this way, the inter-element EMI issue can be fully characterized and considered. However, the extended model implies that the radiation pattern of individual pixel antenna is jointly coded by all the antenna coders due to the EMI issue, making the model complicate and intractable and increasing the complexity for antenna coding optimization.
VI-C2 Spherical Wave
When the spacing is too large, it makes the MIMO pixel antenna system very large, so the assumption of plane wave that we use becomes inaccurate. In this work, we assume that pixel antenna element spacing is not too large, as mentioned above, so that the plane wave assumption is valid. Otherwise, the effect of spherical wave should be incorporated in beamspace channel model for MIMO pixel antenna systems.
VI-D Pixel Antenna System Prototyping
Various pixel antennas have been designed, prototyped, and experimented. For example, in [8] and [10], the experiments for the pixel antenna prototypes have verified that the pixel antennas can flexibly reconfigure the radiation pattern toward different directions. Considering a line-of-sight (LOS) channel scenario, such pixel antennas can be directly used to generate an adaptive beam toward the receiver to enhance the system performance such as channel gain. Thus, the experiments provided in [8] and [10] have preliminarily verified the advantages of pixel antenna system in a simple LOS scenario. However, the previous work [8] and [10] were investigated at the level of antenna hardware design and limited to LOS scenario. To overcome this limitation, in this work we propose antenna coding which considers the impact on wireless communication with multipath fading at the system level to enhance the performance. Thus, in future research, the proposed pixel antenna system can be prototyped similarly to [8] and [10], but needs more efforts in the signal processing to control the pixel antenna with the proposed antenna coding technique.
VII Conclusions
In this paper, we propose a novel technique denoted antenna coding empowered by pixel antennas. Pixel antennas are a flexible reconfigurable antenna technology based on discretizing a continuous radiation surface into small elements termed pixels. The characteristics of the pixel antennas, such as radiation patterns can be controlled by adjusting the connections between adjacent pixels through switches, which can be leveraged to enhance wireless communication systems.
We first derive a physical and EM based communication model for pixel antennas using microwave multiport network theory and beamspace channel representation. With the communication model, we consider the SISO pixel antenna system and optimize antenna coding using the SEBO algorithm to maximize channel gain. To lower computational complexity, we propose a codebook design for antenna coding based on the generalized Lloyd algorithm. We also analyze the average channel gain of SISO pixel antenna systems and derive that the upper bound of average channel gain is given by the EADoF of the pixel antennas. In addition, we jointly optimize the antenna coding and transmit signal covariance matrix to maximize the channel capacity in MIMO pixel antenna systems.
We evaluate the performance of SISO and MIMO pixel antenna systems compared to conventional SISO and MIMO systems. It is shown that using pixel antennas can enhance the average channel gain by up to 5.4 times and the channel capacity by up to 3.1 times, demonstrating the significant potential of pixel antennas as a new dimension to the design and optimization of wireless communication systems.